# KEN KANGDA WREN1

### v7

@@ A Properties of Inferential Hypothesis Testing
    Standard Continuous-Time Inference. Given setup (Section 1.1), the log-LR process for testing time-homogeneous Wiener data of drift $\mathbf{1}_{+}(B)r^{Z}$ and volatility $\sigma^{Z}$ has the form (Peskir and Shirayev (2006)): in terms of signal-to-noise $\sigma^{l}=\frac{r^{Z}}{\sigma^{Z}}$,
@@
    with $\{w_{\tau}\}$ being some standard Wiener process. Time-dependence obtains via an absolutely continuous clock-change $t\mapsto\tau(t)$: image $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is Wiener, with continuous time-varying parameters (e.g. Kallsen (2006)). It and the resulting belief process by Bayes’ Rule (3) read: in terms of a standard Wiener process $\{w_{t}\}$ in $t$-time,
@@
    Non-Resolution and Novikov’s Condition. For inference based on equivalent Wiener-measure pair ${\mathbf{W}}_{+}\sim\mathbf{W}_{-}$, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d{\mathbf{W}}_{+}|_{t}(\cdot)}{d\mathbf{W}_{-}|_{t}(\cdot)}<\infty$ and so a finite total variance for the log-LR process (A.2) almost surely. Risk-resolution (Item-5, Section 1.1) thus equates to the LR process, the exponentiation of (A.2), or equivalently (A.3), violating Novikov’s Condition, and vice versa.
@@
    Regular Inference. It resolves $B$-outcomes predictably and so continuously. With i.i.d data risk-resolution occurs only at $T_{\mathcal{B}}=\infty$. Finite-time resolution can be modelled via a clock-change that brings resolution forward. Note that log-LR processes (A.2) diverge as $t\to T_{\mathcal{B}}\leq\infty$ while the associated beliefs (A.3) stay finite and continuous. For assets with horizon or maturity $T<T_{\mathcal{B}}$ or with $T_{\mathcal{B}}\leq T<\infty$ but with model-risk $B$ not $\mathcal{F}^{Z}_{T}$-measurable, our model-risk pricing applies all the same on $0,T_{\mathcal{B}}\wedge T)$ regardless. In the rest of the cases, where model-risk $B$ is $\mathcal{F}^{Z}_{T}$-measurable and hence replicable/hedgeable, its ’economic pricing’ (Remark 1) becomes irrelevant anyway.
@@ B PROOF OF PROPOSITION 1
Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d\mathbf{W}^{T}_{-,t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-3&5, Section 1.1), we have:
@@
with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ (Item-2, Section 1.1) and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), with $\sigma^{lD}_{t}$ likewise from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
@@
With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$, we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$, and the following, by (8-11) and (19),
@@
Consider first the case of trivial $\{D_{t}\}$, so that the drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$ satisfying $d\hat{W}^{Z}_{B,t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d\hat{RP}^{B}_{t}$, $t<T$.
@@
By Remark 12, we have $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side vanishes on its own, meaning:
@@
This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
@@
Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}RP^{B}(t)$ that stems from $B$-sure risk-pricing:
@@ C THE MOMENTUM AND LOW-RISK FORMULAE
For the assets of Section 3.2-3.3, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
@@
The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices (37) then derive via (3) by change of variable: for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, we have $dl=(v\underline{v})^{-1}dv$, with:
@@
$H_{\Pi\pm}:=\log\circ\text{ }O_{f}[\underline{\Pi}^{1}_{0\pm}]=H_{\mathbf{p}}+\log(\rho K^{\pm 1}),$ (C.3)
$H_{\mathbf{p}}:=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}],$ (C.4)
@@ C.1 Tracking Inferential Progress and Bias Dominance
Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome. We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is ’non-extreme’, to exclude easy inference or profits.
@@ C.2 The Window of Opportunity
Consider profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ and $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ (see (43)), with $t$ fixed inside the window above and $u>0$ large enough for the second event to be outside the window. Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$. Thus, in-window events dominate, with a log-LR of $\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, $t$ fixed.
@@ C.3 Momentum and Volatility Mixtures
$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
@@
Note the source of the conditionalities above, model-risk pricing $K$. For $K=1$, unpriced model-risk, conditional mixes equal the unconditional. For $K\in(1,2)$, competitively priced model-risk, conditional mixes are perturbations of the unconditional.
@@
Both mix-ratios (C.11-C.12) are monotone declining in $v$, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$. That is, for $\rho\gg 1$, in the window (C.8) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level. Mix reverts to its unconditional level as $t\to\infty$ when model-risk resolves. Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
@@ C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
Further, given (C.10) and under condition (C.8), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$. That is, the leading order solution (43) is not affected by these contributions whenever $t_{\rho}/t\gg 1$, i.e. while bias dominates.
@@
Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are invariant under $v\leftrightarrow\underline{v}$ switching (see (40) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming peak-volatility as the location of peak-reward and ’high risk high reward’ under classical conditions.

### v8

@@ A Properties of Inferential Hypothesis Testing
    Standard Continuous-Time Inference. Given setup (Section 1.1), the log-LR process for testing time-homogeneous Wiener data of drift $\mathbf{1}_{+}(B)r^{Z}$ and volatility $\sigma^{Z}$ has the following form (Peskir and Shirayev (2006)): in terms of signal-to-noise $\sigma^{l}=\frac{r^{Z}}{\sigma^{Z}}$,
@@
    with $\{w_{\tau}\}$ being a standard Wiener process. Time-dependence obtains via an absolutely continuous clock-change $t\mapsto\tau(t)$: the image $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is Wiener, with continuous time-varying parameters (e.g. Kallsen (2006)). It and the resulting belief process by Bayes’ Rule (4) then read: in terms of a standard Wiener process $\{w_{t}\}$ in $t$-time,
@@
    Non-Resolution and Novikov’s Condition. For inference using equivalent Wiener-measure pair ${\mathbf{W}}_{+}\sim\mathbf{W}_{-}$, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d{\mathbf{W}}_{+}|_{t}(\cdot)}{d\mathbf{W}_{-}|_{t}(\cdot)}<\infty$ and so a finite total variance for the log-LR process (A.2) and non-resolution. Risk-resolution (Item-6, Section 1.1) thus equates to the LR process, the exponentiation of (A.2), or equivalently (A.3), violating Novikov’s Condition, and vice versa.
@@
    Regular Inference. Uncertainty resolves predictably and so continuously under regular inference. With i.i.d data resolution occurs at $T_{\mathcal{B}}=\infty$. It can be brought forward under a clock-change that transforms processes on $(0,\infty)$ to those on some $(0,T_{\mathcal{B}}\leq\infty)$. Note that log-LR processes (A.2) diverge as $t\to T_{\mathcal{B}}$ while the associated beliefs (A.3) stay finite and continuous. For assets with horizon or maturity $T<T_{\mathcal{B}}$, or with $T_{\mathcal{B}}\leq T<\infty$ but a model-risk $B$ that is not $\mathcal{F}^{Z}_{T}$-measurable, our model-risk pricing theory applies all the same on $0,T_{\mathcal{B}}\wedge T)$. In all other cases, where model-risk $B$ is $\mathcal{F}^{Z}_{T}$-measurable and hence replicable/hedgeable, arbitrage pricing applies.
@@ B PROOF OF PROPOSITION 1
Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d\mathbf{W}^{T}_{-,t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{D}_{t}\}$ being independent standard Wiener processes (Item-4&6, Section 1.1), we have:
@@
with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ (Item-3, Section 1.1) and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), with $\sigma^{lD}_{t}$ likewise from $B$-conditionally independent data $\{D_{t}(B)\}$, so $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
@@
With viable model-risk only pricing, under a RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$, we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$, and the following, by (8-11) and (19),
@@
By Remark 12, $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift requires each side above to vanish on its own, meaning:
@@
It is solved by $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure of the form $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
@@
Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}RP^{B}(t)$ due to $B$-sure risk-pricing:
@@ C THE MOMENTUM AND LOW-RISK FORMULAE
For the assets of Section 3.2-3.3, all the distributions required derive from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
@@
The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices (37) then derive via (4) by change of variable: for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, we have $dl=(v\underline{v})^{-1}dv$, since:
@@
$H_{\Pi\pm}:=\log\circ\text{ }O_{f}[\underline{\Pi}^{1}_{0\pm}]=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}]+\log(\rho K^{\pm 1})=:H_{\mathbf{p}}+\log(\rho K^{\pm 1}),$ (C.3)
@@ C.1 Tracking Inferential Progress and Bias Dominance
Bias thus creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome. We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is ’non-extreme’, to exclude easy inference or profits.
@@ C.2 The Window of Opportunity
Take profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ and $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ (see (43)), with $t$ fixed inside the window above and $u>0$ large enough for the second event to be outside the window. Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$. Thus, in-window events dominate, with a log-LR of $\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, $t$ fixed.
@@ C.3 Momentum and Volatility Mixtures
$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}K^{-(t_{\mathbf{p}}/t)-(-1)^{B}};$ (C.10)
@@
The source of the conditionalities above is model-risk pricing $K$: if $K=1$ (unpriced model-risk), conditional mixes equal the unconditional; if $K\in(1,2)$ (competitively priced model-risk), conditional mixes are perturbations of the unconditional.
@@
Both mix-ratios (C.10-C.11) are monotone declining in $v$, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$. That is, for $\rho\gg 1$, in the window (C.7) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level. Mix reverts to its unconditional level as $t\to\infty$ when model-risk resolves. Any uneven unconditional background merely introduces a constant factor in (C.10-C.11); all above statements remain valid.
@@ C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
Further, given (C.9) and under condition (C.7), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$. That is, the leading order solution (43) is not affected by these contributions under $t_{\rho}/t\gg 1$ (i.e. bias domination).
@@
Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are invariant under $v\leftrightarrow\underline{v}$ switching (see (40) and (C.9-C.11)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming peak-volatility as the location of peak-reward and ’high risk high reward’ under classical conditions.